OpenStudy (anonymous):

(1/(e^x))(sqrt(1+(e^(-x^2))))dx? <-- how?

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty} (1/e^x)\sqrt(1+(1/e^(x^2)))dx <--more specifically\]

OpenStudy (anonymous):

if I do: u = 1/(e^x) du = -e^(-x) I would eliminate the first e^-x giving me: -integral of sqrt(1 + u^2)du but this one would make it messier because of tan^2theta. I checked wolfram and it gave me e^(-x) sqrt(e^(-x^2)+1) but i dunno how this is.

OpenStudy (anonymous):

\(\left(\dfrac1{e^x}\right)^2\neq\dfrac1{e^{-x^2}}\)

OpenStudy (anonymous):

Give us the correct problem please.

OpenStudy (anonymous):

ok sorry. that would be 1/(e^(x^2))

OpenStudy (raden):

retype ur equation from int (........./......) dx

OpenStudy (anonymous):

int ( (1/e^x) (sqrt( 1 + ( 1/ (e^(x^2) ) ) ) dx

OpenStudy (anonymous):

is this any better?

OpenStudy (anonymous):

If it's really \(e^{x^2}\), good luck.

OpenStudy (anonymous):

Yeah, I dunno how to do e^x^2. If nothing else, ill just copy paste what wolfram said and hope the prof thinks i know how to do it.

OpenStudy (anonymous):

Do you mean \(e^{x^2}\) or \((e^x)^2\)?

OpenStudy (anonymous):

(1/(e^x))^2

OpenStudy (anonymous):

aw. sorry. it was originally (1/(e^x))^2 so that would have given me (1/e^2x) I was wrong. aw.

OpenStudy (anonymous):

you know what? just nevermind this. thanks though. ill just skip this for now.